OpenStudy (anonymous):

x^2+y^2-2x+6y+3=0 this equation is equivalent to .... 1. (x-1)^2 +(y+3)^2=-3 2. (x-1)^2+(y+3)^2=7 3. (x+1)^2+(y+3)^2=7 4. (x+1)^2+(y+3)^3=10 help please. explain steps please

OpenStudy (anonymous):

u complete squares, ans is 2.

OpenStudy (anonymous):

can you show me how you did it?

OpenStudy (anonymous):

do you know how to complete a square?

OpenStudy (anonymous):

group the terms like this \[x^2-2x+y^2+6y=-3\]

OpenStudy (anonymous):

then for the x think "half of -2 is -1, and \[(-1)^2=1\]" also think "half of 6 is 3, and \[3^2=9\] " then write \[ (x-1)^2+(y+3)^2=-3+1+9\]